Angelina wants to add 60% antifreeze solution to her car's radiator. About how much 60% antifreeze solution can she make by mixing 1 gallon of pure antifreeze with water?
Answer by ptaylor(2198)

Let x=amount of water needed
Now we know that the amount of pure antifreeze before the water is added (1 gal) has to equal the amount of pure antifreeze after the water is added (0.60(1+x)), so:
1=0.60(1+x)
1=0.60+0.60x subtract 0.60 from each side
1-0.60=0.60-0.60+0.60x collect like terms
0.40=0.60x divide each side by 0.60
x=2/3 gal------------------amt of water needed
1=0.60(1 2/3)=0.60*(5/3)
1=1
Hope this helps---ptaylor